OpenStudy (anonymous):

Use the definition of the derivative to find the derivative of each function with respect to x. f(x)=x^2 - x - 5 Can anyone explain how to do it?

OpenStudy (anonymous):

Do you knw the power rule?

OpenStudy (aum):

\[\large \frac{d}{dx}x^n = nx^{n-1} \\ \large f(x) = x^2 - x - 5 \\ \large f'(x) = \frac{d}{dx}x^2 - \frac{d}{dx}x - \frac{d}{dx}5 = ? \]

OpenStudy (anonymous):

The power rule lets use x^2

OpenStudy (anonymous):

\[x ^{2}= 2x.........3x^3 = 9x^2.........x ^{-4 }=-4x ^{-5}\]

OpenStudy (anonymous):

In the power rule to find the derivative you take the \[x ^{n}=n x ^{n-1}\]

OpenStudy (anonymous):

Do you see this?

OpenStudy (anonymous):

Yes Thank you!

OpenStudy (anonymous):

Ok so for your equation you need the derivative for each value

OpenStudy (aum):

\[ \large \frac{d}{dx}x^n = nx^{n-1} \\ \large f(x) = x^2 - x - 5 \\ \large \frac{d}{dx}f(x) = \frac{d}{dx}x^2 - \frac{d}{dx}x - \frac{d}{dx}5 \\ \large \frac{d}{dx}f(x) =~~~2x~~~ -~~~~ 1 ~~~- ~~~~0 \\ \large \frac{d}{dx}f(x) = 2x - 1 \]